OpenStudy (anonymous):

Choose the correct statement below regarding heat and temperature. The temperature of an object is related to its kinetic energy - increasing the temperature increases the mass in the kinetic energy equation. Heat is energy that is transferred between objects. Only kinetic energy can be transformed into heat energy. When a hot object is placed in contact with a cold object, it loses heat until it reaches the temperature of that cold object.

OpenStudy (anonymous):

y do u hate me

OpenStudy (anonymous):

What? who says I hate you.

OpenStudy (anonymous):

u

OpenStudy (anonymous):

the last one's true...

OpenStudy (anonymous):

You sure? This is a quiz so I can't get it wrong.

OpenStudy (anonymous):

I love you AngelDreamGurl <3 How can someone hate an angel, tell me now?

OpenStudy (anonymous):

But the last statement does make sense,

OpenStudy (anonymous):

glad we agree :) and yeah - i'm pretty sure... -> http://en.wikipedia.org/wiki/Conduction_%28heat%29

OpenStudy (anonymous):

That was wrong. It's the second one

OpenStudy (anonymous):

:'(

OpenStudy (anonymous):

Do you know this one? Which statement is true? Molecules in gases possess kinetic energy since they are in constant motion, while molecules in liquids and solids are not in constant motion and hence possess no kinetic energy. Molecules in gases and liquids possess kinetic energy since they are in constant motion, while molecules in solids are not in constant motion and hence possess no kinetic energy. Polyatomic molecules possess kinetic energy in the liquid and gaseous states since the atoms can move about in the molecule even if the molecule cannot move. Since solids are rigid, their molecules do not possess kinetic energy unless the solid is melted. Molecules in gases, liquids and solids possess kinetic energy since they are in constant motion.

OpenStudy (anonymous):

mah - flutter this quiz! The last statement is as true as the second! (sry - i thought 'heat' was a synonym for 'thermal energy')... It's just guessing what the guy phrasing the questions was thinking .. and apparently he wanted you to know, that the energy is not rly 'lost' but transferred. But how the **** should we was thinking about the word 'lose' in exactly this way..? argh :-/

OpenStudy (anonymous):

And about the second question: I _swear_ the last one is true!

OpenStudy (anonymous):

Yep... :)

OpenStudy (anonymous):

When pure sodium hydroxide is dissolved in water, heat is evolved. In a laboratory experiment to measure the molar heat of solution of sodium hydroxide, the following procedure was followed. To a calorimeter containing 3.00 × 102 g of water at 20.00 °C, 10.65 g of NaOH, also at 20.00 °C was added. The temperature change, which was monitored by a digital thermometer with negligible heat capacity, stopped when it reached a temperature of 28.50 °C. If the specific heat of the mixture is 4.184 J g-1°C1, and the small heat capacity of the calorimeter is ignored, what is the heat evolved, per mole of sodium hydroxide?

OpenStudy (anonymous):

Ur still typing. Appreciate the help :)

OpenStudy (anonymous):

ahm... so you got 306g water + 10.65g NaOH = 316.65g of mixture @ 20°C and then later @ 28.5°C...\[\Delta T = 8.5°\] the "specific heat capacity" (hope that's the correct name) c is given with 4.184 J g^-11 °C^-1 (should be in J kg^-1 K^-1 but whatever) \[\Delta Q = c*m*\Delta T = 4.184 * 316.65 * 8.5 = 11261.3 J\] so about 11.3 kJ for that 10.65g of NaOH. NaOH should be 23+16+1 = 40 g/mol, so we are looking at 10.65/40 = 0.27 mol And therefore, 1 mol of NaOH would give you about 11.3kJ/0.26mol = 42.3 kJ/mol of heat. yeah I am ^^ hope, it's correct - better calculate it again for yourself, cause apparently i'm having a bad day ;)

OpenStudy (anonymous):

These are the choices I have. -41.5 kJ -45.5 kJ -90.5 kJ -37.4 kJ -153 kJ

OpenStudy (anonymous):

argh.. so.. let's plug in the exact numbers and see what we get ^^

OpenStudy (anonymous):

\[(4.184*316.65*8.5) / (10.65/40) = 42296.11492957746\]

OpenStudy (anonymous):

That would round up to be?

OpenStudy (anonymous):

it's 42.3 again - and i honestly have no clue what the guy makin the quiz was thinking, cause that's definitely the correct answer ( i think ^^)... Cant find an error on my side, so - it's your choice - guess what he was doing :P (we are closer to 41.5 - but I wouldnt bet on it - maybe he just divided by 0.25.. you'd already be at 45 with that...)

OpenStudy (anonymous):

The professor is a flutterin idiot. All the quizzes have been like this.

OpenStudy (anonymous):

So 42.3?

OpenStudy (anonymous):

yeah - but I could've made a mistake, too... Best thing to do, would be, u go over it step by step again and double-check everything.. hopefully you find the error, if there is one.

OpenStudy (anonymous):

All right. Do you know this? For a chemical reaction taking place at constant pressure in which all reactants and products are solids or liquids, ΔE ≈ qp. Is it True or False?

OpenStudy (anonymous):

sry, i dont know - what's 'q' ? and what's 'p' ?

OpenStudy (anonymous):

q is the change in heat